Status Report

NASA ISS On-Orbit Status 5 July 2008

By SpaceRef Editor
July 5, 2008
Filed under , , ,
NASA ISS On-Orbit Status 5 July 2008

All ISS systems continue to function nominally, except those noted previously or below. Saturday – off duty for CDR Volkov, FE-1 Kononenko & FE-2 Chamitoff.

Crew Sleep Cycle: Wake/sleep cycle remains right-shifted (5:30am – 9:00pm EDT).

Greg Chamitoff had his second session with the NASA/JSC experiment NUTRITION w/Repository, the FD30 session for blood collection only, for which he had to forego exercising and food intake for eight hours. Later today, the FE-2 will also set up the equipment for the 24-hour urine collections which start with the first void early tomorrow morning and continue through Monday morning. [After performing self-phlebotomy, i.e., drawing blood samples (from an arm vein), the samples were first allowed to coagulate in the Repository for 20-30 minutes, then spun in the HRF RC (Human Research Facility/Refrigerated Centrifuge) and finally placed in MELFI (Minus-Eighty Laboratory Freezer for ISS). No thruster activity was allowed during the blood drawing. The RC was later powered off after a temperature reset to limit wear on the compressor, and cleaned.]

The crew performed the regular weekly three-hour task of thorough station cleaning. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the house cleaning, the crew also conducted regular maintenance inspection & cleaning on fan grilles in the FGB (TsV2), SM (VPkhO, VPrK, FS5, FS6 & FS9), DC1 (V3) screens and PF1/PF2 dust filters, as well as in the COL (Columbus Orbital Laboratory).

In addition, Volkov temporarily powered down the POTOK air filtration system for the periodic cleaning of its pre-filter, using the vacuum cleaner with narrow-slit nozzle attachment.

Earlier, Sergey serviced the Russian BMP (Harmful Impurities Removal System), starting the "bake-out"-to-vacuum cycle on absorbent bed #2 of the regenerable dual-channel filtration system. The regen process will be terminated tonight at ~5:15pm EDT. Filter bed #1 regeneration was performed yesterday. [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP’s regeneration cycle, normally done every 20 days, is currently performed four times more frequently to remove any lingering Freon-218 from the cabin atmosphere (last time done: 6/19&20).]

Later today, FE-1 Kononenko is scheduled for regular maintenance work on the powered-down Russian SRVK-2M water condensate processor, first replacing its BKO multifiltration column unit with a new BKO spare, then also removing & replacing its life-expired FGS gas-liquid mixture filter with a new one. Both old units are to be discarded

Gregory completed today’s routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables.

For the CDR, it was time again for the recharging of the Motorola Iridium-9505A satellite phone brought up on Soyuz 16S, a monthly routine job and his third time. [After retrieving it from its location in the TMA-12/16S descent module (BO), Sergey was to initiate the recharging of its lithium-ion battery, monitoring the process every 10-15 minutes as it takes place. Upon completion at ~4:50pm, the phone will be returned inside its SSSP Iridium kit and stowed back in the BO’s operational data files (ODF) container. The satphone accompanies returning ISS crews on Soyuz reentry & landing for contingency communications with SAR (Search-and-Rescue) personnel after touchdown (e.g., after an “undershoot” ballistic reentry, as happened during the recent 15S return). The Russian-developed procedure for the monthly recharging has been approved jointly by safety officials. During the procedure, the phone is left in its fire-protective fluoroplastic bag with open flap. The Iridium 9505A satphone uses the Iridium constellation of low-Earth orbit satellites to relay the landed Soyuz capsule’s GPS (Global Positioning System) coordinates to helicopter-borne recovery crews. The older Iridium-9505 phones were first put onboard Soyuz in August 2003. The newer 9505A phone, currently in use, delivers 30 hours of standby time and three hours of talk, up from 20 and two hours, respectively, on the older units.]

The three crewmembers conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR), RED resistive exercise device (CDR, FE-1, FE-2), and VELO bike with bungee cord load trainer (CDR, FE-1). Later tonight, Gregory will transfer the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

The Russian crewmembers had their weekly PFCs (Private Family Conferences) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), Sergey at ~9:45am EDT, Oleg at ~11:30am.

At ~7:50am, as part of his FD30 NUTRITION w/Repository session today, Gregory held a teleconference with ground specialists at JSC, including the Principal Investigator, Dr. Scott Smith.

At ~8:05am, the three crewmembers tagged up for 90 min with EVA representatives of star City’s Cosmonaut Corps to go over the two upcoming Russian spacewalks in careful detail, via S-band during several comm windows.

Weekly Science Update (Expedition Seventeen — Week 11)

3-D SPACE : Second session has been successfully performed by Greg Chamitoff on 6/30, “much faster than the first session”.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module. Last memory card exchange was performed on 7/1.

ANITA (Analyzing Interferometer for Ambient Air): Continuing.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Samples returned on 1J.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): The experiment platform continues to operate nominally. A DHPU (Data Handling Processing Unit) software patch is currently being developed to fix the link error issues encountered with DEBIE-2 and FIPEX instruments. – DEBIE-2: Link error still in work. However a work-around allows for regular science data acquisition using an on-board IOP (Instrument Operations Procedure), but not yet in conjunction with any FIPEX IOP – a final software patch, which will enhance the present instrument capabilities, is under finalization on ground; – DOSTEL: On-going science acquisition; – EuTEMP: Currently inactive as planned; – EVC: Several successful ground commanded runs. Some nice pictures of Earth at sunset and sunset could be obtained. In full Sun illumination, the EVC sensor seems to saturate and consequently the picture quality is not optimum; – EXPOSE: On-going science acquisition; – FIPEX: Science script was running nominally until 6/28, when a sudden sensor switch off was encountered. It is a known problem, under assessment. Currently up and running; – MEDET: Some commandings were performed to adjust the frequency of acquisition of the microcalorimeter. Continuously acquiring science data; – PLEGPAY: currently in READY mode, no science data acquisition on-going; – TRIBOLAB: The Ball-Bearing experiment 1/2/3 which is running since 6/16 has been paused during the EUTEF power-down on 6/21 and has been resumed since. On 6/25, the experiment went suddenly to Non-Nominal Mode, after some higher than expected currents have been measured with the motor. On 6/30, the instrument was put in Thermal Stabilization Mode, and the Ball-Bearing experiment 1/2/3 was restarted on 7/1. This experiment run is expected to last until Week #14.

FSL (Fluid Science Laboratory): FSL is back and running. Further FSL commissioning activities remain to be done.

GEOFLOW: A long and complex mechanical procedure which aimed at replacing the lamp of the CEM (Central Experiment Module)-Upper Drawer has been successfully completed by Greg on 7/1. The next day, ground teams in Europe activated the FSL Rack and successfully completed a significant number of the check-out steps. That means that FSL is back and running. The start of GEOFLOW is pending further FSL remaining commissioning activities (FSL RIC and FSL VMU software upgrades, check of FSL optical modes). In the near future, the crew will swap the optical target, stand-by until further optical modes checks, and then will be able to insert the GEOFLOW Experiment Container. “…we are a significant step closer to GEOFLOW start. Thanks a lot for the great job on-orbit!”

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Planned.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer, ESA): Measurements are currently planned before and after the Russian EVA-20 on 7/10 for the two Russian cosmonauts.


PADLES (Passive Dosimeter for Lifescience Experiment in Space): In progress.

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SHERE (Shear History Extensional Rheology Experiment): SHERE was successfully installed into the MSG (Microgravity Science Glovebox) on 6/29.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): In progress.

SOLAR (Solar Monitoring Observatory): LAN (Local Areas Network) connectivity with SOLAR has been lost since 6/25. SOLAR is currently powered via the feeder #2 (survival) only and all instruments are off. Troubleshooting plan has been worked out this week and the next steps would require crew support to diagnose the connectivity between Columbus and SOLAR, making use of the ESA Multi-Purpose Laptop. Unfortunately most of the on-going Sun observation period data acquisition has been lost, which ends on 7/7. – SOVIM: Instrument switched off pending above issue; – SOLSPEC: Instrument switched off pending above issue; – SOLACES: Instrument switched off pending above issue.

SOLO: Planned.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The four ECs (Experiment Containers) of Rotor B are yet to be retrieved by the crew from the blocked Rotor B. BIOLAB could not be activated from ground yet, pending resolution of the Smoke Detector issue. A troubleshooting plan has been developed to recover all the BIOLAB functionalities. The remaining WAICO-#1 ECs will be disposed and replaced by six Reference ECs on Rotor B.

CEO (Crew Earth Observations): Continuing.

No CEO (Crew Earth Observations) photo targets uplinked for today.

CEO photography can be studied at this “Gateway” website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

Significant Events Ahead (all dates Eastern Time, some changes possible.):
07/10/08 — Russian EVA-20A (2:18pm)
07/15/08 — Russian EVA-20 (1:14pm)
07/23/08 — ATV1 reboost (tent.)
09/05/08 — ATV1 undocking, from SM aft port (loiter until ~9/25 for nighttime reentry/observation)
09/09/08 — Progress M-64/29P undocking, from FGB nadir (may move to 8/30)
09/10/08 — Progress M-65/30P launch
09/12/08 — Progress M-65/30P docking (SM aft port)
10/01/08 — NASA 50 Years
10/08/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
10/11/08 — Progress M-65/30P undocking (from SM aft port)
10/12/08 — Soyuz TMA-13/17S launch
10/14/08 — Soyuz TMA-13/17S docking (FGB nadir port)
10/23/08 — Soyuz TMA-12/16S undocking (DC1 nadir)
11/10/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC
11/12/08 — STS-126/Endeavour/ULF2 docking
11/20/08 — ISS 10 Years
11/26/08 — Progress M-66/31P launch
11/28/08 — Progress M-66/31P docking
02/10/09 — Progress M-67/32P launch
02/12/09 — Progress M-67/32P docking
1QTR CY09 — STS-119/Discovery/15A launch – S6 truss segment
03/25/09 — Soyuz TMA-14/18S launch
2QTR CY09 — STS-127/2J/A launch – JEM EF, ELM-ES, ICC-VLD
3QTR CY09 — STS-128/17A/Atlantis – MPLM(P), last crew rotation
05/27/09 — Six-person crew on ISS (following Soyuz 19S docking, May ’09)
3QTR CY09 — STS-129/ULF3/Discovery – ELC1, ELC2
4QTR CY09 — STS-130/20A/Endeavour – Node-3 + Cupola
1QTR CY10 — STS-131/19A/Atlantis – MPLM(P)
1QTR CY10 — STS-132/ULF4/Discovery – ICC-VLD, MRM1 (contingency)
2QTR CY10 — STS-133/ULF5/Endeavour – ELC3, ELC4 (contingency).

SpaceRef staff editor.